QUESTION 8 8.1 Let L1 and L2 be lines defined by x = w0 + su, s ∈ R and y = w1 + tv, t ∈ R, respectively. Show that L1and L2 are parallel if and only if u = kv for some k ∈ R. (8) 8.2 Find the plane that passes through the point (2, 4, −3) and is parallel to the plane −2x + 4y − 5z + 6 = 0. (4) 8.3 Find the line that passes through the point (2, 5, 3) and is perpendicular to the plane 2x − 3y + 4z + 7 = 0. (4) 8.4 Find an equation of the plane passing through the point(−2, 3, 4) and is perpendicular to the line passing through the points (4, −2, 5) and (0, 2, 4). (4)
1
Expert's answer
2020-05-06T19:33:18-0400

8.8 Let "L_1" be defined by "\vec{x} = \vec{w_0} + s\vec{u}, s \in \R" and "L_2" defined by "\vec{y} = \vec{w_1} + t\vec{v}, t \in \R." We may rewrite these equations in a parametric form

"\begin{cases}x_1=w_{0,1}+su_1,\\x_2=w_{0,2}+su_2,\end{cases} 
 \begin{cases}y_1=w_{1,1}+tv_1,\\y_2=w_{1,2}+tv_2\end{cases} ."

We should show that "\exists k\in \R: \, \vec{u}=k\vec{v}", that is,

"u_1 = kv_1, u_2=kv_2."


Let us determine the slope of the lines.

"k_1 = \dfrac{x_2-w_{0,2}}{x_1-w_{0,1}} = \dfrac{su_2}{su_1} = \dfrac{u_2}{u_1},"

"k_2 = \dfrac{y_2-w_{1,2}}{y_1-w_{1,1}} = \dfrac{tv_2}{tv_1} = \dfrac{v_2}{v_1}."

If the two lines are parallel, then "k_1=k_2" , therefore

"\dfrac{u_2}{u_1} = \dfrac{v_2}{v_1}."

This formula means that vectors "\vec{u}" and "\vec{v}" are proportional, therefore "\exists k\in \R: \, \vec{u}=k\vec{v}."


8.2 Let the equation of plane be "A x + B y + C z + D = 0". Two planes are parallel if

"\dfrac{A_1}{A_2} = \dfrac{B_1}{B_2} = \dfrac{C_1}{C_2} = k."

So we get

"\dfrac{A}{-2} = \dfrac{B}{4} = \dfrac{C}{-5} = k."

Let "k=1" , because rescale the equation to get "k( A x + B y + C z + D )= 0." So "A=-2, \; B=4,\; C=-5."

The plane passes through the point "(2, 4, \u22123)" , so we obtain

"-2\cdot2+4\cdot4 -5\cdot (-3)+D = 0,"

therefore "D=-27."

The equation of the plane is "-2x+4y-5z-27=0."


8.3 Let the equation of the line be "\vec{r}=\vec{r_0} + t\vec{a},"

where "\vec{r_0} = (2,5,3)" is radius vector of point "(2,5,3)." The plane defined by "Ax+By+Cz+D=0" and line are perpendicular if "\vec{a}" is parallel to "\vec{v}=(A,B,C),"

so we get

"\dfrac{a_x}{A}=\dfrac{a_y}{B}=\dfrac{a_z}{C}" ,

therefore

"\dfrac{a_x}{2}=\dfrac{a_y}{-3}=\dfrac{a_z}{4}" .

We may choose "a_x = 2,\; a_y=-3, \; a_z = 4." So the equation of the line may be rewritten in parametric form as

"\begin{cases} x = x_0 + a_xt,\\ y=y_0 + a_yt, \\ z=z_0 + a_zt \end{cases}" and "\begin{cases} x = 2 +2t,\\ y=5 -3t, \\ z=3 + 4t .\end{cases}"


8.4 Let the equation of plane be "A x + B y + C z + D = 0" .

First we determine the equation of the line passing through two given points:

"\dfrac{x-x_1}{x_2-x_1} = \dfrac{y-y_1}{y_2-y_1}=\dfrac{z-z_1}{z_2-z_1}."

Therefore

"\dfrac{x-4}{0-4} = \dfrac{y+2}{2+2}=\dfrac{z-5}{4-5}, \;\; 
\dfrac{x-4}{-4} = \dfrac{y+2}{4}=\dfrac{z-5}{-1}, \;\;" so the line is parallel to vector "(-4,4,-1)."

As the plane and line are perpendicular. we get

"\dfrac{A}{-4}=\dfrac{B}{4} = \dfrac{C}{-1}."

We may choose "A=-4, \; B=4,\; C=-1." We know that plane passes through "(-2,3,4)," so

"-2A+3B+4C+D=0, \;\; -2\cdot(-4) + 3\cdot4 +4\cdot(-1) + D = 0,"

therefore "D=-16."

So the equation of the plane is "-4x+4y-z-16=0."